Calculate the molarity of a solution when 600 grams of NaC2H3O2 are added to 2 L of solution


1
Expert's answer
2021-05-19T02:27:38-0400

NaC2H3O2 have molar mass =


23 + 2×12 + 3 × 1 + 2×16

= 23 + 24 + 3 + 32

= 82 g / mol .


moles of NaC2H3O2

= 600 / 82

= 7.32 mol.


Molarity = moles / Volume .

= 7.32 / 2

= 3.66